OpenStudy (anonymous):

For an angle Θ with the point (16, −30) on its terminating side, what is the value of cosine?

OpenStudy (anonymous):

\(16\) over the hypotenuse of the triangle with one side \(16\) and the other \(30\)

OpenStudy (anonymous):

i.e. \[\frac{16}{\sqrt{16^2+30^2}}\] as a start

OpenStudy (anonymous):

You need to solve for c. a2+b2=c2

OpenStudy (anonymous):

@PrincessHush nice picture!

OpenStudy (anonymous):

Thanks! (: @satellite73

Directrix (directrix):

Doesn't theta in this case have one side on the x-axis?